# T2VSafetyBench: Evaluating the Safety of Text-to-Video Generative Models

### v1

@@ Abstract
The recent development of Sora leads to a new era in text-to-video (T2V) generation. Along with this comes the rising concern about its security risks. The generated videos may contain illegal or unethical content, and there is a lack of comprehensive quantitative understanding of their safety, posing a challenge to their reliability and practical deployment. Previous evaluations primarily focus on the quality of video generation. While some evaluations of text-to-image models have considered safety, they cover fewer aspects and do not address the unique temporal risk inherent in video generation. To bridge this research gap, we introduce T2VSafetyBench, a new benchmark designed for conducting safety-critical assessments of text-to-video models. We define 12 critical aspects of video generation safety and construct a malicious prompt dataset using LLMs and jailbreaking prompt attacks. Based on our evaluation results, we draw several important findings, including: 1) no single model excels in all aspects, with different models showing various strengths; 2) the correlation between GPT-4 assessments and manual reviews is generally high; 3) there is a trade-off between the usability and safety of text-to-video generative models. This indicates that as the field of video generation rapidly advances, safety risks are set to surge, highlighting the urgency of prioritizing video safety. We hope that T2VSafetyBench can provide insights for better understanding the safety of video generation in the era of generative AI.
@@ 1 Introduction
To bridge this research gap, in this work we propose T2VSafetyBench, a new benchmark for evaluating the safety of text-to-video models. By examining the usage policies of OpenAI, LLaMa-2, and Anthropic and surveying dozens of AI safety practitioners, we identify 12 critical aspects of video generation safety: Pornography, Borderline Pornography, Violence, Gore, Public Figures, Discrimination, Political Sensitivity, Illegal Activities, Disturbing Content, Misinformation and Falsehoods, Copyright and Trademark Infringement, and Temporal Risk. To evaluate these aspects, we generate multiple malicious text prompts using LLMs and various jailbreaking prompt attacks against diffusion models [38, 41, 27], followed by manual screening and fine-tuning. For the generated videos, we capture a frame per second and use these multi-frame images along with the manually designed prompts to assess safety via GPT-4. Given that automated metrics might not accurately reflect human judgement on safety, we also conduct manual assessments and calculate the correlation between GPT-4 assessments and human evaluations.
@@ 2 Related work
Text-to-video generation and evaluation. Text-to-Video (T2V) generation using latent diffusion model has taken a significant leap in the past two years [37, 16, 6, 10, 1, 29, 17, 3, 44]. Make-A-Video [37] and Imagen-Video [16] train a cascaded video diffusion model, making researchers see the hope of purely AI-generated videos. LVDM [14], Align Your latent [7] and MagicVideo [51] extend latent text-to-image model to the video domain through additional temporal attention or transformer layer. Text2Video-Zero [19] enables zero-shot video generation from textual prompts, while Stable Video Diffusion [6] can achieve multi-view synthesis from a single image. VideoPoet [20] leverages autoregressive language model to perform multitasking across various video-centric inputs and outputs. Commercial text-to-video models like Gen2 [10] and Pika [1] also play a pivotal role in this field. The recent phenomenal Sora [29] adopts DiT [30] as backbone to generate high-fidelity 1-minute video from text and strictly adhere to user instructions. However, Sora is close-sourced currently thus we adopt one of its alternatives named Open-Sora [17]. Several benchmarks [23, 18, 25] evaluate video generation quality, including aspects such as text alignment, motion quality, and temporal consistency. Nevertheless, text-to-video models face significant safety challenges, as generated videos may contain illegal or unethical content, synthetic identities, misinformation, and potential infringements of copyrights or privacy [24]. Current benchmarks have not adequately addressed these safety concerns.
Safety benchmark for generative models. Generative large models, such as ChatGPT [35] and Stable Diffusion [36], can produce unsafe content [55, 8, 34], raising widespread concern. PromptBench [53] initially investigates the robustness of large language models (LLMs) against adversarial prompts. DecodingTrust [42] evaluates several perspectives of trustworthiness in GPT models. A series of studies [50, 5, 46, 32, 22, 26, 28, 49] further assesses the safety risks associated with LLMs and multimodal large language models (MLLMs). Additionally, several works [21, 33, 47] have evaluated the safety of text-to-image models. HEIM [21] provides a holistic evaluation of text-to-image models, including evaluations of toxicity. Qu et al. [33] explores the factors contributing to the generation of hateful memes, Yang et al. [47] proposes a new method to generate adversarial prompts. However, these studies have limited their focus to insufficient aspects such as pornography, violence, gore, hate, and politics, neglecting other critical safety aspects. They also fail to consider the unique temporal risks associated with video. Our work addresses these gaps by conducting a comprehensive safety assessment of video generation models across 12 crucial dimensions.
@@ 3.1 Aspects
Finally, Temporal Risk is a unique safety dimension associated with videos compared to images. The temporal risk dimension focuses on the risks associated with video content over time, where individual frames may appear harmless but the entire sequence can convey inappropriate content or themes through interaction, continuity, or dynamic changes between frames. For example, as shown in Figure 2, a sequence of seemingly benign screens in Times Square, through specific ordering and timing, may subtly reveal NSFW content. This dimension requires models to consider not only the superficial safety of individual frames but also to analyze and understand the context and potential implications of the entire sequence.
@@ 3.2 Dataset construction
We focus on evaluating text-to-video models that take textual prompts as inputs to generate corresponding videos. We employ OpenAI’s GPT-4 [2] to generate multiple malicious text prompts for each aspect and manually screen and fine-tune these prompts. Furthermore, we implement various methods of jailbreaking prompt attacks against diffusion models [38, 41, 27] to more effectively gather malicious prompts capable of generating inappropriate videos for a more thorough evaluation.
@@ 3.2.1 Dataset construction based on LLMs
Initially, we generate multiple malicious text prompts for each aspect using GPT-4 [2]. The detailed instructions provided to GPT-4 are shown in Table 3. Although we intentionally emphasize the diversity of test data in our prompt instructions, LLMs still tend to increase the probability of repeating previous sentences, resulting in a self-reinforcement effect [45]. We mitigate this by manually removing prompts that convey meanings similar to existing malicious prompts to ensure dataset diversity. Additionally, to ensure the quality of the generated prompts, we rigorously review and fine-tune harmful prompts to maintain consistency with the definitions of their respective aspects.
@@ 3.2.2 Dataset construction based on prompt attacks
where $g$ is the generator, $\mathcal{L}_{harm}$ quantifies the harmfulness of the generated $g(\hat{P})$, $\mathcal{L}_{sim}$ evaluates the similarity between $P$ and $\hat{P}$, $\mathcal{L}_{tox}$ measures the manifest toxicity of $\hat{P}$. Specifically, BSPA uses a black-box LLM (e.g., GPT-4) to produce $\hat{P}$ from $P$, then employs the pre-defined text and image filters (e.g., detoxify [13] and stable diffusion safety checker [5]) to calculate the toxicity of $\hat{P}$. After that, BSPA conducts a text retriever to find the most related sensitive word (i.e., the most harmful words while maintaining high similarity to original prompt $P$) to update $\hat{P}$. These attack methods enable more effective finding of high-quality harmful prompts for thorough evaluation.
@@ 3.3 Evaluation protocols
Evaluating the safety of text-to-video generation presents challenges as it requires comprehensive and fine-grained cross-modal understanding. To enhance the assessment across 12 aspects, we employ the multimodal large language model GPT-4 to determine the safety of the generated videos. For each generated video, we capture an image frame per second and present these multi-frame images alongside prompts to GPT-4. These prompts are specifically manually designed to guide the model in classifying whether a given video is safe or not. As depicted in Figure 3, to align with our safety standards, we prompt potential harmful aspects. Considering that the outputs from GPT-4 are sometimes lengthy sentences that do not categorically classify as safe or unsafe, we require the output in the format "ANS: X, Yes: a%, No: b%", to measure the degree to which the generated videos are considered inappropriate. If GPT-4 does not respond in this format, such as by describing the video content without making a judgment, the query is repeated until the response is in the correct format, which is generally effective.
Given that automated metrics may not accurately reflect human judgments of safety [21], we also incorporate a manual evaluation process to compare the effectiveness of GPT-4’s assessments with human judgments. This manual review involves cross-validation by four individuals with expertise in artificial intelligence. We observe that the correlation between GPT-4’s evaluations and manual reviews is typically high. This finding suggests that leveraging GPT-4 on a large scale for evaluation is reasonable in our context.
@@ 4 Benchmarking results
In this section, we thoroughly evaluate the safety of T2V models based on T2VSafetyBench. Subsequently, we conduct an empirical analysis on the evaluation results, delving deeper into the findings. We respectively generate 400 malicious text prompts for 12 aspects and evaluate popular text-to-video models including Pika [1], Gen2 [10], Stable Video Diffusion [6], and Open-Sora [17] (as an alternative to Sora). Both GPT-4 and human assessments are employed. A video is deemed NSFW (Not Safe For Work) if its unsafety score exceeds 0.5. In Table 2, we report the NSFW rate across different models under various aspects as assessed by both GPT-4 and humans, along with the correlation coefficient (CC) between these two evaluations. A higher NSFW rate indicates a higher safety risk. The visualization of the results is shown in Figure 4. Below, we first detail the results of different aspects in Sec. 4.1, then describe the main findings in Sec. 4.2.
@@ 4.1 Perspectives from different aspects
Pornography 23.0% 31.1% 0.0% 0.0% 0.0% 1.6% 45.9% 45.9% 0.844
Borderline Pornography 55.0% 50.0% 30.0% 25.0% 0.0% 5.0% 20.0% 25.0% 0.871
Violence 55.0% 65.0% 60.0% 50.0% 55.0% 55.0% 95.0% 95.0% 0.832
Gore 88.5% 84.6% 0.0% 7.7% 15.4% 19.2% 57.7% 61.5% 0.863
Public Figures 97.3% 97.3% 100.0% 100.0% 81.1% 81.1% 97.3% 86.5% 0.820
Discrimination 24.0% 32.0% 8.0% 16.0% 36.0% 40.0% 20.0% 28.0% 0.831
Political Sensitivity 13.3% 16.7% 60.0% 66.7% 66.7% 46.7% 33.3% 26.7% 0.711
Illegal Activities 50.0% 55.3% 47.4% 47.4% 65.8% 65.8% 47.4% 42.1% 0.683
Disturbing Content 68.0% 92.0% 20.0% 28.0% 48.0% 56.0% 92.0% 76.0% 0.589
Misinformation 66.7% 70.0% 50.0% 56.7% 73.3% 73.3% 80.0% 73.3% 0.752
Copyright and Trademark 19.0% 14.3% 71.4% 66.7% 76.2% 85.7% 38.1% 33.3% 0.881
Temporal Risk 75.0% 85.0% 5.0% 0.0% 0.0% 0.0% 5.0% 5.0% 0.887
NSFW Average 52.9% 57.8% 37.6% 38.7% 43.1% 44.1% 52.2% 51.2% 0.826

### v2

@@ Abstract
The recent development of Sora leads to a new era in text-to-video (T2V) generation. Along with this comes the rising concern about its security risks. The generated videos may contain illegal or unethical content, and there is a lack of comprehensive quantitative understanding of their safety, posing a challenge to their reliability and practical deployment. Previous evaluations primarily focus on the quality of video generation. While some evaluations of text-to-image models have considered safety, they cover fewer aspects and do not address the unique temporal risk inherent in video generation. To bridge this research gap, we introduce T2VSafetyBench, a new benchmark designed for conducting safety-critical assessments of text-to-video models. We define 12 critical aspects of video generation safety and construct a malicious prompt dataset including real-world prompts, LLM-generated prompts and jailbreak attack-based prompts. Based on our evaluation results, we draw several important findings, including: 1) no single model excels in all aspects, with different models showing various strengths; 2) the correlation between GPT-4 assessments and manual reviews is generally high; 3) there is a trade-off between the usability and safety of text-to-video generative models. This indicates that as the field of video generation rapidly advances, safety risks are set to surge, highlighting the urgency of prioritizing video safety. We hope that T2VSafetyBench can provide insights for better understanding the safety of video generation in the era of generative AI.
@@ 1 Introduction
To bridge this research gap, in this work we propose T2VSafetyBench, a new benchmark for evaluating the safety of text-to-video models. By examining the usage policies of OpenAI, LLaMa-2, and Anthropic and surveying dozens of AI safety practitioners, we identify 12 critical aspects of video generation safety: Pornography, Borderline Pornography, Violence, Gore, Public Figures, Discrimination, Political Sensitivity, Illegal Activities, Disturbing Content, Misinformation and Falsehoods, Copyright and Trademark Infringement, and Temporal Risk. To evaluate these aspects, we build a malicious text prompt dataset containing real-world prompts collected from VidProM [53], generated prompts by GPT-4, and various jailbreak attack-based prompts against diffusion models [48, 51, 33], followed by manual screening and fine-tuning. For the generated videos, we capture a frame per second and use these multi-frame images along with the manually designed prompts to assess safety via GPT-4. Given that automated metrics might not accurately reflect human judgement on safety, we also conduct manual assessments and calculate the correlation between GPT-4 assessments and human evaluations.
@@
Ethical Considerations. Our work involves exposure of human reviewers to upsetting content; therefore, we implement a series of safety measures for human evaluators to mitigate potential risks. The key measure includes informing volunteers in advance about the possibility of encountering distressing content, providing examples, and clarifying that they can withdraw from the study at any time without penalty if they feel uncomfortable. Additional safety measures are detailed in Appendix A. We have discussed our procedures and the details of human evaluations with the Institutional Review Board (IRB) and obtained an exempt decision. Additionally, we discuss in Appendix A the potential bias that may arise due to the high cultural specificity of human reviewers and the possibility that prompts used in these benchmarks might over-correct and censor certain kinds of information, potentially causing discrimination. Furthermore, we will carefully consider how to share our dataset responsibly. For instance, to avoid adverse societal impacts, we will release the jailbreak prompts dataset only upon request and for research purposes.
@@ 2 Related work
Text-to-video generation and evaluation. Text-to-Video (T2V) generation using latent diffusion model has taken a significant leap in the past two years [46, 18, 6, 10, 1, 35, 19, 3, 54]. Make-A-Video [46] and Imagen-Video [18] train a cascaded video diffusion model, making researchers see the hope of purely AI-generated videos. LVDM [16], Align Your latent [7] and MagicVideo [64] extend latent text-to-image model to the video domain through additional temporal attention or transformer layer. Text2Video-Zero [23] enables zero-shot video generation from textual prompts, while Stable Video Diffusion [6] can achieve multi-view synthesis from a single image. VideoPoet [24] leverages autoregressive language model to perform multitasking across various video-centric inputs and outputs. Commercial text-to-video models like Gen2 [10] and Pika [1] also play a pivotal role in this field. The recent phenomenal Sora [35] adopts DiT [37] as backbone to generate high-fidelity 1-minute video from text and strictly adhere to user instructions. However, Sora is close-sourced currently thus we adopt one of its alternatives named Open-Sora [19]. Several benchmarks [29, 21, 31, 20, 47] evaluate generation quality, including aspects such as text alignment, motion quality, and temporal consistency. Nevertheless, text-to-video models face significant safety challenges, as generated videos may contain illegal or unethical content, synthetic identities, misinformation, and potential infringements of copyrights or privacy [30]. Current benchmarks have not adequately addressed these safety concerns.
Safety benchmark for generative models. Generative large models, such as ChatGPT [42] and Stable Diffusion [44], can produce unsafe content [68, 8, 41], raising widespread concern. PromptBench [66] initially investigates the robustness of large language models (LLMs) against adversarial prompts. DecodingTrust [52] evaluates several perspectives of trustworthiness in GPT models. A series of studies [62, 5, 57, 39, 28, 32, 34, 60, 61] further assesses the safety risks associated with LLMs and multimodal large language models (MLLMs). Additionally, several works [27, 40, 58] have evaluated the safety of text-to-image models. HEIM [27] provides a holistic evaluation of text-to-image models, including evaluations of toxicity. Qu et al. [40] explores the factors contributing to the generation of hateful memes, Yang et al. [58] proposes a new method to generate adversarial prompts. However, these studies have limited their focus to insufficient aspects such as pornography, violence, gore, hate, and politics, neglecting other critical safety aspects. They also fail to consider the unique temporal risks associated with video. Our work addresses these gaps by conducting a comprehensive safety assessment of video generation models across 12 crucial dimensions.
@@ 3.1 Aspects
Finally, Temporal Risk is a unique safety dimension associated with videos compared to images. The temporal risk dimension focuses on the risks associated with video content over time, where individual frames may appear harmless but the entire sequence can convey inappropriate content or themes through interaction, continuity, or dynamic changes between frames. Temporal Risk can be classified into three types: (1) Sequential Action Risk (where individual frames extracted from actions are benign, but the continuous sequence of actions is not safe for work), (2) Dynamic Variation Risk (where unsafe content is revealed through dynamically transforming the target into another object over time), (3) Coherent Contextual Risk (where unsafe content emerges from the temporal linking of presented elements, such as harmful subtitles). The evaluation criteria for each category are consistent: “the entire sequence presents NSFW content through continuity between frames”. For example, as shown in Figure 2, a sequence of seemingly benign screens in Times Square, through specific ordering and timing, may subtly reveal NSFW content. This dimension requires models to consider not only the superficial safety of individual frames but also to analyze and understand the context and potential implications of the entire sequence.
@@ 3.2 Dataset construction
To evaluate these aspects, we construct our malicious text prompt dataset based on real-world prompts collected from VidProM, generated prompts by GPT-4, and jailbreak attack-based prompts. Firstly, we collect NSFW prompts from the VidProM [53], which contains text-to-video prompts from real users. Secondly, we employ OpenAI’s GPT-4 [2] to generate multiple malicious text prompts for each aspect and manually screen and fine-tune these prompts. Thirdly, we implement various methods of jailbreaking prompt attacks against diffusion models [48, 51, 33] to more effectively gather malicious prompts capable of generating inappropriate videos for a more thorough evaluation. Ultimately, the T2VSafetyBench prompt dataset comprises 4,400 prompts.
3.2.1 Dataset construction based on VidProM
First, we collect real-world prompts from VidProM [53], which is a large-scale dataset comprising 1.67 million unique text-to-video prompts from real users. Based on the NSFW probabilities assigned by the state-of-the-art NSFW model Detoxify [15], we select prompts with an NSFW probability exceeding 0.8. We review and curate these selected prompts, incorporating 2,325 into T2VSafetyBench. Compared to generating malicious prompts directly with LLMs, selecting from VidProM enhances the data sources and better reflects the prompts in the real-world.
@@ 3.2.2 Dataset construction based on LLMs
To further expand and diversify the dataset, we generate multiple malicious text prompts for each aspect using GPT-4 [2]. The detailed instructions provided to GPT-4 are shown in Table 3. Although we intentionally emphasize the multiformity of test data in our prompt instructions, LLMs still tend to increase the probability of repeating previous sentences, resulting in a self-reinforcement effect [56]. We mitigate this by manually removing prompts that convey meanings similar to existing malicious prompts to ensure dataset variety. However, this is still insufficient. To further increase the diversity of prompts, we also employ the Self-Instruct [55] framework. We construct the seed set using previous data, which includes prompts from VidProM and prompts generated by GPT-4 in this section, thereby incorporating both real-world and LLM-generated prompts. Subsequently, we apply Self-Instruct, leveraging the seed set to guide GPT-4 in generating a broader and more diverse range of prompts. Additionally, to ensure the quality of the generated prompts, we rigorously review and fine-tune harmful prompts to maintain consistency with the definitions of their respective aspects. Ultimately, GPT-4 generates a total of 1230 prompts.
@@ 3.2.3 Dataset construction based on prompt attacks
where $g$ is the generator, $\mathcal{L}_{harm}$ quantifies the harmfulness of the generated $g(\hat{P})$, $\mathcal{L}_{sim}$ evaluates the similarity between $P$ and $\hat{P}$, $\mathcal{L}_{tox}$ measures the manifest toxicity of $\hat{P}$. Specifically, BSPA uses a black-box LLM (e.g., GPT-4) to produce $\hat{P}$ from $P$, then employs the pre-defined text and image filters (e.g., detoxify [14] and stable diffusion safety checker [5]) to calculate the toxicity of $\hat{P}$. After that, BSPA conducts a text retriever to find the most related sensitive word (i.e., the most harmful words while maintaining high similarity to original prompt $P$) to update $\hat{P}$. These attack methods enable more effective finding of high-quality harmful prompts for thorough evaluation. Ultimately, jailbreaking prompt attack methods generate a total of 845 prompts.
@@ 3.3 Evaluation protocols
Evaluating the safety of text-to-video generation presents challenges as it requires comprehensive and fine-grained cross-modal understanding. To enhance the assessment across 12 aspects, we employ the multimodal large language model GPT-4 to determine the safety of the generated videos. For each generated video, we sample an image frame per second and present these multi-frame images alongside prompts to GPT-4. Given that existing models produce videos with only gradual and subtle changes between frames, a one-frame-per-second sampling frequency is sufficient to detect harmful content. Actually, the high correlation observed between GPT-4’s evaluations and human evaluations in the experiment also validates that the one-frame-per-second sampling frequency for GPT-4 assessment is effective in this context. This also represents one of the key findings and contributions of our paper. These prompts for GPT-4 are specifically manually designed to guide the model in classifying whether a given video is safe or not. As depicted in Figure 3, to align with our safety standards, we prompt potential harmful aspects. Considering that the outputs from GPT-4 are sometimes lengthy sentences that do not categorically classify as safe or unsafe, we require the output in the format "ANS: X, Yes: a%, No: b%", to measure the degree to which the generated videos are considered inappropriate. If GPT-4 does not respond in this format, such as by describing the video content without making a judgment, the query is repeated until the response is in the correct format, which is generally effective.
Given that automated metrics may not accurately reflect human judgments of safety [27], we also incorporate a manual evaluation process to compare the effectiveness of GPT-4’s assessments with human judgments. We recruit 60 volunteers for the assessment, all of whom must be at least 18 years old, in good physical and mental health, and free from conditions such as heart disease or vasovagal syncope. Before the assessment, we present definitions and examples of each type of safety risk to the volunteers. Volunteers view full videos on 22-24 inch monitors. Volunteers are given a 10-minute break after every 20 minutes of review to ensure psychological comfort and sustained attention. Each video receives evaluations from at least three volunteers. Following the initial evaluations, we conducted a secondary round of cross-validation. We observe that the correlation between GPT-4’s evaluations and manual reviews is typically high. This finding suggests that leveraging GPT-4 on a large scale for evaluation is reasonable in our context.
@@ 4 Benchmarking results
In this section, we thoroughly evaluate the safety of T2V models based on T2VSafetyBench.  [^1] The code for T2VSafetyBench is available at the following link (https://anonymous.4open.science/r/T2VSafetyBench_Code-1763/) for reproduction purposes. Contact {miaoyibo@amss.ac.cn}. Subsequently, we conduct an empirical analysis on the evaluation results, delving deeper into the findings. We preliminarily evaluate popular text-to-video models including Pika [1], Gen2 [10], Stable Video Diffusion [6], and Open-Sora 1.1 [19] (as an alternative to Sora). For each prompt in our dataset, we generate four videos given a T2V model. Videos generated by Open-sora have a duration of 16 seconds with 8 frames per second. Videos generated by Pika have a duration of 3 seconds, while videos generated by Gen2 and SVD have a duration of 4 seconds, all with 24 frames per second. Both GPT-4 and human assessments are employed. A video is deemed NSFW (Not Safe For Work) if its unsafety score exceeds 0.5. In Table 2, we report the NSFW rate across different models under various aspects as assessed by both GPT-4 and humans, along with the correlation coefficient (CC) between these two evaluations. A higher NSFW rate indicates a higher safety risk. The visualization of the results is shown in Figure 4. Below, we first detail the results of different aspects in Sec. 4.1, then describe the main findings in Sec. 4.2.
@@ 4.1 Perspectives from different aspects
Pornography 22.3% 30.4% 0.4% 0.9% 0.1% 1.6% 49.2% 49.8% 0.845
Borderline Pornography 54.5% 51.3% 36.5% 31.1% 1.3% 5.7% 19.7% 24.1% 0.867
Violence 54.3% 65.6% 63.6% 55.2% 56.8% 56.2% 95.9% 95.2% 0.832
Gore 95.2% 91.1% 0.0% 4.0% 19.4% 24.3% 57.4% 61.8% 0.856
Public Figures 97.0% 96.4% 100.0% 100.0% 84.6% 82.5% 97.3% 87.2% 0.818
Discrimination 20.2% 28.7% 8.8% 16.2% 39.7% 44.7% 22.0% 30.7% 0.829
Political Sensitivity 10.6% 14.3% 59.3% 67.2% 70.2% 49.6% 31.8% 24.5% 0.709
Illegal Activities 51.1% 58.3% 47.8% 49.9% 66.3% 66.5% 50.7% 47.5% 0.682
Disturbing Content 73.4% 97.8% 26.0% 35.9% 53.6% 63.0% 93.0% 83.2% 0.602
Misinformation 67.8% 72.8% 47.6% 54.4% 77.0% 78.0% 81.3% 76.6% 0.755
Copyright and Trademark 13.1% 10.3% 76.4% 71.6% 74.2% 85.5% 44.5% 41.8% 0.880
Temporal Risk 81.3% 90.6% 10.1% 4.3% 2.7% 3.5% 3.7% 3.2% 0.889
NSFW Average 53.4% 59.0% 39.7% 40.9% 45.5% 46.8% 53.9% 52.1% 0.826
Pornography. Pika and Open-Sora exhibit a high NSFW rate due to lack of ability to detect and prevent the generation of sexual content. In contrast, Gen2 and SVD demonstrate robust defenses against sexual content. Nearly all malicious prompts are detected by their built-in safety filters, preventing the generation of videos. This disparity stems from Open-Sora lacking detection capability for NSFW content, while Pika only implements a preliminary detector for input text. On the other hand, Gen2 and SVD feature post-generation detectors for the videos themselves, enabling effective identification and rejection of any generated videos containing sexual content.
Borderline Pornography. Pika maintains a relatively high NSFW rate, posing the highest safety risk. In contrast, Gen2 shows a reduction in safety for pornography, while SVD still effectively mitigates such risks. It could be argued that SVD is nearly impeccable in filtering sexual content. Open-Sora exhibits a significant reduction in its NSFW rate for pornography, due to its weaker comprehension ability. Specifically, Open-Sora fails to capture the subtly implicit sexual content in borderline pornography, thereby ensuring the videos it generates are invariably safe.
Violence. All of four video generation models demonstrate elevated NSFW rates in the context of violence. Pika, Gen2, and SVD exhibit NSFW rates between 50% and 65%, while Open-Sora reaches nearly 100%. These findings indicate that current models are suboptimal and lack safety in handling violent content, necessitating further research to address these challenges.
Gore. In the aspect of Gore, Gen2 achieves optimal performance with a NSFW rate close to 0%. SVD closely follows, effectively rejecting the majority of video generations containing gory scenes. Similar to the situation with pornography, due to the lack of detection capability for output videos, both Pika and Open-Sora generate a significant number of gory videos.
Public Figure. None of the four text-to-video generative models incorporate specific detection measures for public figures, although SVD does offer limited detection capability. The relatively low Human NSFW rate of 87.2% in Open-Sora stems from a minor proportion of unrecognizable public figures being generated. In contrast, the GPT-4 NSFW rate reaches 97.3%, possibly due to looser detection standards. Overlooking the detection of public figures could lead to the generation of infringing content and pose legal risks, which necessitates further attention.
Discrimination. All of four models exhibit lower NSFW rates, suggesting minimal safety risks. However, this outcome primarily arises because current video generation models generally possess limited capability, typically generating clear and simple actions rather than effectively representing more abstract and complex content such as discrimination. Consequently, a lower NSFW rate does not imply a robust defense mechanism against discrimination. Even for simple discriminatory actions, such as a single gesture, these models struggle to detect and reject the generation of such content.